OpenStudy (donka74):

solve the following system using substitution....gold medal -5x+y=-2 -3x+6y=-12

OpenStudy (mathstudent55):

Here's a way of doing it. Solve the second equation for 6y. Then multiply the entire first equation by 6 and solve it for 6y. Then equate the two expressions in x that 6y is equal to. Solve for x. Then substitute the value of x in the first original equation and solve for y.

OpenStudy (anonymous):

-6(-5x+y)=-12 30x+-6y=75 -3x+6y=12 27x=65 x=2.3 -3(2.3)+6y=12 -7+6y=12 6y=19 y=3.17

OpenStudy (mathstudent55):

@mddrummer 1) you didn't use substitution 2) your solution is not correct